OpenStudy (anonymous):

Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: How many liters of orange juice are present in the can? Show your work.

OpenStudy (anonymous):

@David27

OpenStudy (flvs.net):

Which one do you want to do?

OpenStudy (anonymous):

part A pls :)

OpenStudy (aaronandyson):

Let "x" be the total number of liters of orange juice and water combined. 80% of the mixture is pure orange juice which means 20% of it is water. Therefore, x * 20/100 = 1 is the equation in single variable that can be used to find x. Therefore, x/5 = 1 Therefore, x = 5

OpenStudy (flvs.net):

0.2x = 1 where x is the total number of liters of orange juice and water combined.

OpenStudy (aaronandyson):

So there is a total of 5 liters of orange juice and water mixture. 1 L of it is water. Therefore, the rest of 5-1 = 4 liters is orange juice present in the can.

OpenStudy (anonymous):

ok so Aaron is part A and Flvs is part B right?

OpenStudy (aaronandyson):

Well.I've answered both parts.

OpenStudy (flvs.net):

Part A) 0.2x = 1 where x is the total number of liters of orange juice and water combined. Part B) 0.2x = 1 Solve for x. Orange juice is 80%. The amount of orange juice is: x * 80 / 100. Put x from solving the above equation.

OpenStudy (anonymous):

thnx guys :)

OpenStudy (flvs.net):

Np, @AaronAndyson you get a medal.